# Get a Grip: Reconstructing Hand-Object Stable Grasps in Egocentric Videos

### 25 Dec 2023 v1

@@ 4.4 Optimising Object Relative Pose
$E(P_{o2h},P_{o2\phi},\{R_{\phi}^{n}\},s;\{\theta^{n}\},\{T_{h2c}^{n}\})=\sum\limits_{n=1}^{N}\lambda_{1}E_{mask}^{n}+\lambda_{2}E_{ins}^{n}+\lambda_{3}E_{clo}^{n}$
@@
The term $E_{mask}$ focuses on estimating a reconstruction that best matches the 2D projections of the object masks throughout the sequence. We measure the error via sum of pixel differences:
@@
where $\mathcal{M}_{o}^{n}$ is the object mask and $\Pi(\cdot)$ is the differentiable projection function. Similar to our optimisation for hand vertices using the masks, we wish not to penalise parts of the object occluded by the hand. Accordingly, we define the mask $\mathcal{C}_{o}^{n}$ such that $\mathcal{C}_{o}^{n}$ is 1 for the object and the background, and 0 for the hand.
@@ 1 ${}^{\dagger}$ 11 11 339
ARCTIC-Grasps (ours) 2023
@@
EPIC-Grasps (ours) 2023
@@
Table 1: Dataset Comparison: $*{}^{*}$: cannot be used for hand-object reconstruction as object poses or segments are not provided. ${}^{\dagger}$: subjects in the releasted train/val set. Two subjects are in the held out test set.
Similar to [50, 5, 19], we use the physical heuristics $E_{ins}$, which pushes the object out of the penetrating region against the hand and a balancing loss $E_{clo}$ which pulls the object to touch these contact regions. Please refer to the appendix for more details of the physical heuristics.
The latent axis $P_{o2\phi}$ is initialised at the origin in the object coordinate system and initially points in the $+z$ direction; the $R_{\phi}^{n}$ are initialised to 0. The initialisation of $P_{o2h}$ is described in the implementation details.
4.5 Hand-Object Pose Refinement
In Sec 4.3 the initial hand mesh was refined without the knowledge of the in-hand object. After the optimisation of $T_{o2h}^{n}$, we keep the object relative poses fixed and perform two steps of refinement in an object-aware manner.
Refining hand-object as a single mesh. We glue the per-frame pair of meshes of the hand and the object into a single mesh, i.e. $V_{ho:h}$ as a concatenation of $V_{h:h}$ and $V_{o:h}$. This step refines the hand pose $T_{h2c}^{n}=[R_{h}^{n}|t_{h}^{n}]$ so that the combined hand-object mesh aligns better with the combined hand-object mask (Fig 3 right-most column). We then optimise the sum of pixel differences and the smoothness of the object in the camera coordinate system $V_{o:c}^{n}$:
$E(\{R_{h}^{n}\},\{t_{h}^{n}\})$ $=F_{s}(\{V_{o:c}^{n}\})+\sum_{n=1}^{N}E_{mask}^{n}$ (10)
$E_{mask}^{n}$ $=|\mathcal{C}_{ho}^{n}\otimes(\mathcal{M}_{ho}^{n}-{\Pi(V_{ho:c}^{n})})|^{2}_{2}$ (11)
where $\mathcal{C}_{ho}^{n}$ is 1 for the hand-object and the background, and 0 for other objects; $M_{ho}^{n}$ is the combined hand-object mask.
Refining finger articulation. As the object has been introduced to the hand, we refine finger articulations such that the hand and object are in plausible contact, using physical constraints $E_{ins}$ and $E_{clo}$ introduced in Sec 4.4:
$E(\{\theta^{n}\})=F_{s}(\{\theta^{n}\})+\sum_{n=1}^{N}E_{ins}^{n}+E_{clo}^{n}$ (12)
As summarised in Fig 3, we perform three stages of optimisation. The first stage refines FrankMocap output using ground-truth hand masks. The second optimises the in-hand object pose using known hand poses and ground-truth object masks. The third refines the hand pose and the finger articulation in an object-aware manner.
5 EPIC-Grasps Annotations
With the definition of stable grasp finalised on in-lab videos of the ARCTIC dataset, we now use this to annotate the temporal extents of stable grasps in unscripted egocentric recordings capturing in-the-wild hand-object stable grasps. This offers a dataset distinct from prior works, where frame-level labels [49, 20, 5], 3D supervision [2, 41, 53] or recordings specifically collected to evaluate grasps with no underlying action [17, 26, 6] are typically used. Instead, we aim to find stable grasps from within egocentric videos recorded of daily actions in an unscripted manner.
We next detail our pipeline to annotate EPIC-Grasps:
1. Identifying candidate clips to annotate. The ultimate goal of hand-object reconstruction is to operate for any rigid or dynamic objects, including novel previously unknown objects. However, current approaches for reconstruction of unknown objects are still in their infancy [51, 52, 39, 22, 40]. We thus restrict our first-attempt at reconstruction in-the-wild to objects for which a category CAD model can be acquired. Note that this is distinct from instance CAD models in ARCTIC-Grasps, for example, the general CAD model of a bottle might not exactly match all bottles in daily life.
We exclude very small objects and shortlist 9 object categories frequently used in kitchens: plate, bowl, bottle, cup, mug, can, pan, saucepan, glass [^1] For the object mesh, we made the CAD model in Blender [8] as the common shape of each category..
We begin with the frame-level hand-object interaction annotations from VISOR [12], in which we identified clips where the a hand is in contact with one of the categories above. The annotations also inform us which hand is in contact with the object (i.e. left or right). We then use the fine-grained action labels available from EPIC-KITCHENS [11, 10] to locate the initial temporal extent for actions involving any of these categories. This resulted in 5516 in-action candidate segments.
2. Annotating Stable Grasps. Two annotators were asked to label the start-and-end frames following the stable grasp definition in Sec 3. We discard segments when (i) a segment does not contain any stable grasp (ii) both the hand and object are out-of-view during the sequence (iii) the object does not match the category CAD model specified.
3. 2D Hand and Object Segmentation Annotations. In total, we labelled 2431 video clips of stable gasps from 141 distinct videos in 31 kitchens [11]. For each clip, we have a start and end time of the stable grasp, as well as 319,661 segmentation masks for the hand and the object during the stable grasp from the dense VISOR annotations [12]. Of these, 1446 contain left hand stable grasps and 985 contain right hand stable grasps. Note that the majority of subjects recording the EPIC-KITCHENS dataset are right-handed—which implies the left hand is more frequently used to steady the object during manipulation. Compared to previous efforts [17, 30, 26, 6, 40], most clips in EPIC-Grasps contains both hands in view manipulating the same or different objects.
@@ 5 EPIC-Grasps Annotations
Table 1 compares our dataset with currently available and regularly used datasets for hand-object reconstruction. In the first section, we show datasets collected with 3D Ground-Truth. These datasets are limited in the environmental setup to allow for multi-camera views or motion capture tracking. This limits the appearance diversity and interactions are not functional. In Table 1 we also report the functional intent of various dataset (i.e. object is used with intent). our dataset is the first to capture in-the-wild egocentric videos with functional intent. Of particular note is the HOI4D [30] where diversity is targeted with a large number of indoor rooms. However, this is a constructed setup, where objects are selected from a given pool. In EPIC-Grasps, subjects are recording in their actual (personal) kitchens and objects.
@@ 6 Experiments
The experiment section is structured as follows: in Sec 6.1 and Sec 6.2 we describe the implementation details of our method and the quantitative metrics respectively. Sec 6.3 and Sec 6.4 report results of our method and baselines on ARCTIC-Grasps and EPIC-Grasps, respectively. Finally Sec 6.5 conducts ablation studies on our method.
@@ 6.1 Implementation Details
Finding stable grasps on ARCTIC. We measure the object in-contact area using set of vertices. As individual vertices are sensitive to noise, we partition vertices into fragments using the farthest-point-sampling method, and measure the area using unions of fragments. We set the number of fragments to 0.01 of the number of vertices. As our pipeline assumes rigid object, we exclude segments where the hand is changing the object’s articulation.
Optimisation Pipeline. We follow the convention in HOMan [19] which sparsely and linearly samples 30 frames from the sequence for optimisation and evaluation. We use 50 initialisations of object rotation and 1 global translation for each object [^2] We use prior knowledge of grasps from ground-truth of ARCTIC to find these initialisations. The optimisation process takes 1.5 minutes on a 2080Ti for one 30-frame sequence on average.
@@
Please refer to appendix for further implementation details.
@@ 6.2 Baselines and Quantitative Metrics
We use 4 baselines to compare to:
@@
Table 2: Results on ARCTIC-Grasps. Green shows the best performing method per metric and yellow shows second best.
  • HOMan [19]—a common baseline in other works [33, 52] that progressively optimises the object.
  • Single Frame Single-frame optimisation without utilising the knowledge of stable grasps.
  • Static A version of our method where objects are not allowed any motion when within the stable grasp, minimising the overall rotation and translation of the object.
  • Dynamic A variation of our method where objects are allowed to move freely within 6-DoF.
We propose quantitative metrics to measure the correctness of the predicted object poses as well as correctness of the stable grasp. We utilise our stable grasp annotations so as to correctly assess reconstruction with/without 3D ground truth. Specifically, given the knowledge that stable grasps maintain a consistent contact area between the hand and the object throughout the sequence, we measure this contact area over time. However, a reconstruction can maintain a stable contact area but be completely erroneous. We thus propose to combine a measure of correctness of reconstruction with the stability of the contact area to produce a robust metric. We explain this next.
Average Distance (ADD %). Following [25, 47, 21], we measure the distance of corresponding vertices between GT and predicted object vertices in the hand coordinate, and average this distance over vertices and frames. ADD is 1 for a sequence if the average distance is less than 10% of the object’s diameter, and 0 otherwise.
Average Stable Contact Area at ADD Success (SCA-ADD %). When a pose is considered correct for a sequence, i.e. ADD is 1, we measure the Stable Contact Area across the sequence, defined as the average IOU of surface in contact area between each pair of frames 3. SCA-ADD is set to 0 when ADD is 0. We average over all examples.
Intersection-over-Union (IOU %). We use IOU as a proxy of pose accuracy when 3D GT is not available. We measure the Intersection-over-Union between the ground-truth mask and the rendered mask for the object in camera view. We report average IOU across all frames. Due to occlusion with other components in the scene, only the non-occluded area of the rendered projection is used.
@@
The first 2 metrics, ADD and SCA-ADD, are evaluated when 3D ground truths are available—e.g. ARCTIC-Grasps dataset; and the other two are the corresponding 2D proxy, which we use for EPIC-Grasps.
@@ 6.3 Results on ARCTIC-Grasps
Table 2 compares results on ARCTIC-Grasps dataset using all the metrics. We focus solely on the object stable grasp optimisation in the ARCTIC-Grasps dataset, thus using ground-truth hand mesh and camera parameters. We also use a known scale for the object. We thus do not use the hand refinement stages to report these results.
While Dynamic and 1-DoF both have significantly higher ADD results, 1-DoF maintains the best SCA when grasps are successfully reconstructed (high SCA-ADD). While Static has considerably high SCA, its IOU and ADD are significantly lower than 1-DoF in every category. HOMan only performs best on the “microwave” object sequences. Single Frame performs worst in every case.
@@ 6.4 Results on EPIC-Grasps
Table 3 compares results on EPIC-Grasps dataset using proxy metrics IOU and SCA-IOU. We also report SCA-IOU at IOU threshold 0.6 since the overall IOU for are lower. We evaluate two best variations based on 3D performance on ARCTIC-Grasps: Dynamic and 1-DoF. 1-DoF achieves the best SCA-IOU metric for both thresholds for every object category and overall.
HOMan [19] Dynamic 1-DoF
IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6
bottle 56.7 3.4 5.8 64.8 7.0 20.4 62.7 10.2 33.1
bowl 54.4 1.1 1.6 58.5 7.6 12.4 57.2 12.2 22.2
can 48.3 3.4 5.1 52.4 10.1 15.4 52.4 12.7 24.6
cup 56.9 5.5 7.8 61.4 6.0 24.2 60.3 10.1 36.8
glass 55.4 3.0 4.2 60.2 8.2 21.7 57.5 10.2 29.7
mug 59.4 3.9 6.6 59.0 5.0 18.0 60.1 6.7 28.7
pan 48.4 0.5 1.9 47.0 2.5 6.9 47.5 4.4 13.5
plate 61.1 1.0 1.6 67.1 12.9 19.3 66.6 18.7 30.3
saucepan 51.1 0.4 2.4 57.2 3.4 11.5 55.9 9.0 19.7
All 54.9 1.9 3.3 58.6 7.2 15.5 57.8 10.8 25.2
Table 3: Results on EPIC-Grasps. Green for best method and yellow shows second best.
Figure 4 shows reconstruction results of our 1-DoF approach on the EPIC-Grasps dataset. We showcase three examples per category. From the images, the challenge in EPIC-Grasps dataset is evident with the hand and object occupying a small proportion of the image. In most examples, the palm and fingers are concealed. Importantly, all grasps are functional – e.g. picking a cup from the drainer.
Figure 4: Qualitative Results on EPIC-Grasps (3 examples per category); input frames (top), projected reconstruction results (bottom) and reconstruction in rotated views (right).
Figure 5 shows comparative results. Despite acceptable overlay, HOMan does not maintain the stable grasp and the object is in fact sliding in the hand over time. The Dynamic baseline has significantly lower SCA with the contact area drifting over time. Our proposed 1-DoF method is able to reconstruct the stable grasp as it benefits from the constrained relative pose assumption.
Figure 5: Qualitative comparison between 1-DoF, Dynamic and HOMan [19]. We show input frames (top), grasp reconstruction results and the corresponding SCA.
While results in-the-wild are very promising, our pipeline relies on FrankMocap as a first stage. Despite the robustness incorporated by the multiple-view joint optimisation, our method fails when the majority of hand poses are incorrect (see Fig 6).

### 07 Apr 2024 v2

@@ 3.3 Reconstructing Object Poses in a Stable Grasp
$E(\underbrace{\phi,\{\omega^{n}\},T_{o2h}}_{\{T_{o2h}^{n}\}},s;\{\theta^{n}\},\{T_{h2c}^{n}\})=\sum\limits_{n=1}^{N}\lambda_{1}E_{mask}^{n}+\lambda_{2}E_{push}^{n}+\lambda_{3}E_{pull}^{n}$ (5)
where $\{\theta^{n}\}$ and $\{T_{h2c}^{n}\}$ are sets of outputs from HaMeR and are kept fixed, and $s$ is the scalar object scale to be optimised.
We initialise the latent axis $\phi$ to the object’s z-axis from the CAD model; $\{\omega^{n}\}$ are initialised to zero. We note the initialisations for $T_{o2h}$ in implementation details. We then jointly optimise all parameters across frames, in particular the rotation axis $\phi$ and the per-frame rotation angles $\{\omega^{n}\}$.
@@
We use three terms in the optimisation: $E_{mask}^{n},E_{push}^{n},E_{pull}^{n}$. We describe these next. The main term $E_{mask}$ focuses on estimating a reconstruction that best matches the 2D projections of the object masks throughout the sequence. We measure the error via sum of pixel differences:
@@
where $\mathcal{M}_{o}^{n}$ is the object mask which we use for supervision and $\Pi(\cdot)$ is the differentiable projection function [23]. $\mathcal{C}_{o}^{n}$ is the occlusion-aware mask as in [18, 50] which only computes the error within regions of the object that are not occluded by the hand, set to 1 for the object and the background, and 0 for the hand. This masking is critical to avoid penalising the missing parts of the object mask due to in-hand occlusion.
We also employ two additional terms, used in previous works [46, 4, 18, 31]. We use the physical heuristics $E_{push}$, which pushes the object out of the penetrating region against the hand and a balancing loss $E_{pull}$ which pulls the object to touch these contact regions. For exact calculations of these previously used physical heuristics terms, refer to the appendix.
4 EPIC-Grasps Annotations
With the definition of stable grasp in Sec 3.2, we now use this to annotate the temporal extents of stable grasps in unscripted egocentric recordings capturing in-the-wild hand-object stable grasps. This offers a dataset distinct from prior works, where frame-level labels [45, 19, 4], 3D supervision [2, 39, 49] or recordings specifically collected to evaluate grasps with no underlying action [17, 25, 5] are typically used. Instead, we aim to find stable grasps from within egocentric videos recorded of daily actions in an unscripted manner.
We next detail our pipeline to annotate EPIC-Grasps:
1. Identifying candidate clips to annotate. The ultimate goal of hand-object reconstruction is to operate for any rigid or dynamic objects, including novel classes. However, current approaches for reconstruction of unknown objects are still in their infancy [47, 48, 37, 21, 38] (we show failure cases in appendix). We thus restrict our first-attempt at reconstructing stable grasps in-the-wild to objects of a known category. Note that this is distinct from instance CAD models – the general CAD model of a bottle might not exactly match all bottles in daily life. We exclude very small objects and shortlist 9 object categories frequently used in kitchens: plate, bowl, bottle, cup, mug, can, pan, saucepan, glass [^2] For the object mesh, we made one category CAD model in Blender [7]..
We begin with the frame-level hand-object interaction annotations from VISOR [11], in which we identified clips where the a hand is in contact with one of the categories above. The annotations also inform us which hand is in contact with the object (i.e. left or right). We then use the fine-grained action labels available from EPIC-KITCHENS [10, 9] to locate the initial temporal extent for actions involving any of these categories.
2. Annotating Stable Grasps. Two annotators were asked to label the start-and-end frames following the stable grasp definition. We discard segments when (i) a segment does not contain any stable grasp (ii) both the hand and object are out-of-view during the sequence or (iii) the object does not match the category CAD model specified.
3. 2D Hand and Object Segmentation Annotations. In total, we label 2431 video clips of stable gasps from 141 distinct videos in 31 kitchens [10]. For each clip, we provide a start and end time of the stable grasp, as well as 319,661 segmentation masks for the hand and the object during the stable grasp from the dense VISOR annotations [11]. Of these, 1446 contain left hand stable grasps and 985 contain right hand stable grasps. Note that the majority of subjects recording the EPIC-KITCHENS dataset are right-handed—which implies the left hand is more frequently used to steady the object during manipulation. Compared to previous efforts [17, 28, 25, 5, 38], most clips in EPIC-Grasps contains both hands in view manipulating the same or different objects. We provide the breakdown per category and grasp duration in appendix.
@@ 1 ${}^{\dagger}$ 11 11 339
ARCTIC-Grasps (ours) 2024
@@
EPIC-Grasps (ours) 2024
@@
Table 1: Dataset Comparison: $*{}^{*}$: cannot be used for hand-object reconstruction as object poses or segments are not provided. ${}^{\dagger}$: subjects in the released train/val set.
@@
Tab. 1 compares our dataset with currently available and regularly used datasets for hand-object reconstruction. In the first section, we show datasets collected with 3D ground truth and add the 1303 stable grasp sequences previously noted in Sec. 3.2. We refer to these as ARCTIC-Grasps. We note that previous datasets, including those that requiring 3D ground truth, are limited in the environmental setup to allow for multi-camera views or motion capture tracking. This limits the appearance diversity and interactions are not functional. In Tab. 1 we also report the functional intent of various dataset (i.e. object is used with intent). EPIC-Grasps is the first to capture in-the-wild egocentric videos with functional intent. Of particular note is the HOI4D [28] where diversity is targeted with a large number of indoor rooms. However, this is a constructed setup, where objects are selected from a given pool. In EPIC-Grasps, subjects are recording in their personal kitchens and using their chosen objects as part of longer activities.
@@ 5 Experiments
The experiment section is structured as follows: in Sec. 5.1 and Sec. 5.2 we describe the implementation details of our method and the quantitative metrics respectively. Sec. 5.3 and Sec. 5.4 report results of our method and baselines on ARCTIC-Grasps and EPIC-Grasps, respectively. Finally Sec. 5.5 conducts ablation studies on our method.
@@ 5.1 Implementation Details
We follow the convention in HOMan [18] which sparsely and linearly samples 30 frames from the sequence for optimisation and evaluation. We use 50 initialisations of object rotation and 1 global translation for each object, from ground truth of ARCTIC to find these initialisations, following [18]. The optimisation process takes 1.5 minutes on a 2080Ti for one 30-frame sequence on average.
@@ 5.2 Baselines and Quantitative Metrics
We focus on baselines that do not require training on the dataset in advance. We thus use 4 baselines to compare to:
  • HOMan [18]—a common CAD-based baseline that progressively optimises the object pose relative to the hand. We use the ground truth masks as in our method for fair comparison.
  • Single Frame—We predict the hand and object meshes independently for each frame, without utilising any temporal optimisation or the knowledge of stable grasps.
  • Static—A version of our method where objects are not allowed any motion when within the stable grasp, minimising the overall rotation and translation of the object.
  • Dynamic—A variation of our method where objects are allowed to move freely within 6-DoF.
We note that RHOV [31] does not have released code to be used as a baseline.
We propose quantitative metrics to measure the correctness of the predicted object poses within the stable grasp. Given the knowledge that stable grasps maintain a consistent contact area between the hand and the object throughout the sequence, we measure this contact area over time. However, a reconstruction can maintain a stable contact area but be completely erroneous. We thus propose to combine a measure of correctness of reconstruction with the stability of the contact area to produce a robust metric. We explain this next.
Average Distance (ADD %). Following [24, 43, 20], we measure the distance of corresponding vertices between GT and predicted object vertices in the hand coordinate system, and average this distance over vertices and frames. ADD is 1 for a sequence if the average distance is less than 10% of the object’s diameter, and 0 otherwise.
Average Stable Contact Area at ADD Success (SCA-ADD %). When a pose is considered correct for a sequence, i.e. ADD is 1, we measure the stable contact area across the sequence, defined as the average IOU of in contact area between each pair of frames (Sec. 3.2). SCA-ADD is set to 0 when ADD is set to 0 (average distance below threshold). We average SCA-ADD over all examples.
Intersection-over-Union (IOU %). We use IOU as a proxy of pose accuracy when 3D GT is not available. We measure the Intersection-over-Union between the ground truth mask and the rendered mask for the object in camera view. We report average IOU across all frames. Due to occlusion with other components in the scene, only the non-occluded area of the rendered projection is used.
@@
Table 2: Results on ARCTIC-Grasps. Green shows the best performing method per metric and yellow shows the second best.
Figure 6: Two stable grasp reconstructions by 1-DoF method on ARCTIC-Grasps.
@@
The first two metrics, ADD and SCA-ADD, are evaluated when 3D ground truths are available—e.g. ARCTIC-Grasps dataset; and the other two are the corresponding 2D proxy, which we use for EPIC-Grasps.
@@ 5.3 Results on ARCTIC-Grasps
Tab. 2 compares results on ARCTIC-Grasps, using all the metrics. We here focus on the object metrics, using ground truth hand mesh and a known object scale. Whilst Dynamic and 1-DoF both have significantly higher ADD results, 1-DoF maintains the best SCA when grasps are successfully reconstructed (high SCA-ADD). Static has considerably high SCA, but IOU and ADD are significantly lower than 1-DoF in every category. HOMan only performs best on the “microwave” object sequences. Single Frame performs worst in every case. We show qualitative examples in Fig. 6.
HOMan [18] Dynamic 1-DoF
Category IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6
bottle 56.7 3.4 5.8 75.5 21.8 36.6 72.1 26.1 51.3
bowl 54.4 1.1 1.6 58.0 9.3 19.8 56.2 11.1 25.7
can 48.3 3.4 5.1 56.0 13.7 20.1 54.1 16.5 26.6
cup 56.9 5.5 7.8 67.5 16.5 38.9 65.9 18.4 46.1
glass 55.4 3.0 4.2 65.2 14.8 30.1 62.5 15.7 37.8
mug 59.4 3.9 6.6 63.8 6.3 29.6 62.6 10.8 38.6
pan 48.3 0.5 1.9 48.9 4.3 12.8 49.1 7.0 20.4
plate 61.1 1.0 1.6 68.1 17.6 30.1 65.2 16.3 34.6
saucepan 51.1 0.4 2.4 57.5 5.4 25.7 56.8 8.5 34.6
All 54.9 1.9 3.3 61.7 12.2 25.3 59.9 14.1 32.9
Table 3: Results on EPIC-Grasps. Green for best and yellow shows the second best.
Figure 7: Qualitative comparison between 1-DoF, Dynamic and HOMan [18] on EPIC-Grasps. We show input frames, reconstruction results and the corresponding SCA.
@@ 5.4 Results on EPIC-Grasps
Tab. 3 compares results on EPIC-Grasps dataset using proxy metrics IOU and SCA-IOU at two thresholds (0.8 and 0.6). We compare 1-DoF to the best variation from ARCTIC-Grasps: Dynamic. 1-DoF achieves the best SCA-IOU metric for both thresholds for every object category and overall.
Fig. 7 shows comparative results. Despite acceptable overlay, HOMan does not maintain the stable grasp and the object is in fact sliding in the hand over time. The Dynamic baseline has significantly lower SCA with the contact area drifting over time. Our proposed 1-DoF method is able to reconstruct the stable grasp as it benefits from the constrained relative pose assumption.
Fig. 8 shows reconstruction results of our 1-DoF approach on the EPIC-Grasps dataset. We showcase three examples per category. From the images, the challenge in EPIC-Grasps is evident with the hand and object occupying a small proportion of the image. In most examples, the palm and fingers are concealed. Importantly, all grasps are functional – e.g. picking a cup from the drainer.